# Parametric Operator Inference to Simulate the Purging Process in Semiconductor Manufacturing

### v1

@@ 5 Conclusion
We simulated the purging process in the PECVD chamber using OpInf ROMs, which leverage a data-driven approach to model dynamical systems with complex multiscale and nonlinear behavior. The OpInf framework interpolated between nine ROMs constructed using different argon flow rates at the inlet and outlet pressures in the purging process. This enabled the prediction of the system’s behavior for 16 unseen parameter combinations. The OpInf ROMs maintained high predictive accuracy for the pressure and temperature across the entire spatiotemporal domain. While velocity predictions showed relatively higher errors, particularly near the wafer edge, the ROM captured critical flow dynamics in the wafer center region, which is essential for particle contamination control. In general, the parametric OpInf ROMs showed good predictive accuracy across 25 parameter combinations, with a maximum error of 9.32%. Notably, this approach substantially lowered computational costs, achieving a 142-fold speedup in online computations compared to the FOM CFD simulation. We selected a low training-to-testing ratio to demonstrate the predictive accuracy of the interpolatory OpInf ROM (see Figure 5(a)). However, using a more conventional ratio, e.g., 20 training and five testing datasets, would enhance the results further. This fast and accurate predictive model for the flow field in the PECVD purging process enables the prediction of particle movement within the PECVD chamber, making it a useful tool for particle contamination control. This capability will be particularly crucial in the fast-paced and precision-driven semiconductor manufacturing environment.
Future work could explore developing models that predict particle behavior based on the flow field predictions from parametric OpInf ROMs, which would enable the simulation of particle contamination dynamics. Additionally, matching the numerical predictions of particle behavior with particle measurement data from the PECVD chamber would further improve predictive accuracy.

### v2

@@ 5 Conclusion
We simulated the purging process in the PECVD chamber using OpInf ROMs, which leverage a data-driven approach to model dynamical systems with complex multiscale and nonlinear behavior. The OpInf framework interpolated between nine ROMs constructed using different argon flow rates at the inlet and outlet pressures in the purging process. This enabled the prediction of the system’s behavior for 16 unseen parameter combinations. The OpInf ROMs maintained high predictive accuracy for the pressure and temperature across the entire spatiotemporal domain. While velocity predictions showed relatively higher errors near the wafer edge, the ROM captured critical flow dynamics in the wafer center region, which is sensitive to particle contamination. In general, the parametric OpInf ROMs showed good predictive accuracy across 25 parameter combinations, with a maximum error of 9.32%. This approach substantially lowered computational costs, achieving a 142-fold speedup in online computations compared to the FOM CFD simulation. Overall, this fast and accurate predictive model for the flow field in the PECVD purging process enables the prediction of particle movement within the PECVD chamber, making it a useful tool for particle contamination control. This capability will be particularly crucial in the fast-paced and precision-driven semiconductor manufacturing environment.
There are several interesting directions for future work. First, the performance of the proposed approach may vary depending on the choice of training and testing datasets. We selected a low training-to-testing ratio (9 train/16 test) to demonstrate the predictive accuracy of the ROM, see Fig. 5(a). While adopting a more conventional split (e.g., 20 training and five testing datasets) would likely improve overall accuracy, our goal was to highlight the model’s interpolation performance under limited training. An interesting direction for future work would be to explore the ROM’s extrapolation capability beyond the training domain—for example, in cases where $\mu_{p}<0.80$ or $\mu_{q}<0.56$. Second, we note that the contamination-related analysis presented in this work is inferential, relying on flow-field predictions, rather than explicit particle tracking or contamination metrics. Specifically, the surrogate model is validated through the accuracy of the velocity components most directly governing particle transport—the radial and y-velocity fields in the contamination-critical wafer center region—rather than through direct simulation of particle trajectories. A natural extension would be to couple particle-tracking models with the flow-field predicting ROMs. This integration would enable direct simulation of particle transport and deposition, and comparison with particle measurement data from the PECVD chamber, thereby offering a more comprehensive framework for contamination risk prediction. Lastly, it would also be interesting to investigate whether this approach would work for faster transient phenomena, such as fast atomic layer deposition steps, where the flow dynamics vary significantly on much smaller time scales.